OpenStudy (anonymous):

Solve A = 2p – 3r for r.

OpenStudy (ganpat):

A = 2p - 3r since we are solving for r.. get r on one side and rest all terms on other.. So, 3r = 2p -A r = (2p - A)/ 3 is this the answer ??

OpenStudy (karatechopper):

ok! so we are solving for r! A=2p-3r A-2p=2p-2p-3r A-2p=-3r A-2p/-3=-3r/-3 A-2p/-3=r

OpenStudy (karatechopper):

@Ganpat you forgot that there was a negative sign in front of the 3..

OpenStudy (ganpat):

@karatechopper : we both have same answers.. i have multiplied negative sign inside the brackets..

OpenStudy (anonymous):

THESE ARE THE CHOICES A.) r = A – 2p B.) r = A – 2p + 3 C.) r = negative A plus 2 p, divided by 3 D.) r = negative A plus 3 p, divided by 2

OpenStudy (ganpat):

c. Third one , its the same what i have mentioned..

OpenStudy (anonymous):

Thats what i was going to ask, thank you for clarifying. I appriciate it, thank you once again. & thanks @karatechopper

OpenStudy (ganpat):

w/c..

OpenStudy (karatechopper):

yea i think C